# Retrieval-Based Video Language Model for Efficient Long Video Question Answering

### v1

@@ 3.3 End-to-End Optimization
We train the network using an end-to-end optimization approach with the video instruction data. The image encoder, text encoder, and the LLM are frozen during the training. Only the parameters of the MLP block and projector are optimized. For a video-text pair, besides the cross-entropy loss between the generated prediction and groudtruth answer $\mathcal{L}^{pred}_{i}$, we introduce soft matching (SM) loss $\mathcal{L}_{i}^{\text{SM}}$ to regularize the similarity learning as
@@
$\mathcal{L}_{i}^{\text{SM}}=-\cos(\mathbf{q}_{i},\bar{\mathbf{v}}_{i}),\quad\text{where}\>\bar{\mathbf{v}}_{i}=\frac{\sum_{j=1}^{L_{i}}e^{s_{i}^{j}}\mathbf{v}_{i}^{j}}{\sum_{j=1}^{L_{i}}e^{s_{i}^{j}}}.$ (3)
Here $\bar{\mathbf{v}}_{i}$ is a weighted combination of the $\mathcal{L}_{i}$ chunk features, with weights determined by the similarities between the query and chunk features (see Eq.(2)). In order to maximize the cosine similarity score between query and the combined feature, the MLP block needs to be optimized to result in higher similarity to those question relevant chunks. Better optimized similarity scores lead to better selection of chunks to the LLM and thus superior performance.
The overall loss is as
@@
where $\lambda$ is a hyper-parameter that balances the contribution of the regularization term. We determine the value of $\lambda$ such that both loss terms are on the same order of magnitude. We set $\lambda=10$ in our experiments.
4 Experiments
4.1 Implementation Details
@@
Following Video-ChatGPT (Maaz et al., 2023), we use the Language-aligned Large Vision Assistant (LLaVA) (Liu et al., 2023) as our base model. We utilize the pre-trained CLIP ViT-L/14 (Radford et al., 2021) as our image encoder and extract the feature from the second-to-last layer as the $h\times w$ vision tokens of a frame. We use the fine-tuned Vicuna (7B) from LLaVA as our LLM. For the text encoder, we use the pre-trained CLIP ViT-L/14 text encoder and extract the class token feature of the penultimate layer. The number of neurons for the two fully connected layers of the MLP block is 1024 and 1024, respectively. The number of neurons for the projector is 4096.
@@
We only fine-tune the MLP block and the projector while keeping the image encoder, the text encoder and the LLM frozen. We fine-tune the model for 3 epochs using video instruction data, with a learning rate of 2e-5 and a batch size of 40. Training our model takes about 24h on an A100 80GB GPU. We set $K$ to 5, resulting in $5\times(64+4)=340$ vision tokens as the input to the LLM. This is comparable to the number of vision tokens used in Video-ChatGPT (356 vision tokens).
@@
Table 1: Information of the evaluation datasets, including the number of videos, the number of QA pairs, and the average video duration.
Dataset #Vid #QA Duration (sec.)
WildQA 261 652 71.2
QaEgo4D 166 1854 495.1
lifeQA 49 372 74.0
Social-IQ 2.0 144 876 60.0
4.2 Datasets
@@
We use Video Instruction Data collected by Maaz et al. (2023) for video instruction tuning. This dataset contains about 100k question and answer pairs based on the Activitynet dataset (average duration 180 seconds). The 100k QA pairs include various types of questions, including but not limited to describing the general content of the video, questions related to appearance, movement, trajectory, reasoning, and some tasks that require imagination.
We evaluate the generalization performance of our framework on four video datasets: WildQA (Castro et al., 2022), QaEgo4D (Bärmann & Waibel, 2022), lifeQA (Castro et al., 2020), and Social-IQ 2.0 (Wilf et al., 2023). The average duration of these datasets is larger than one minute. The average duration of QaEgo4D is more than eight minutes. Please see Table 1 for specific information about each dataset. Note that long video datasets are rare. Many popular video QA datasets are not suited for our study, where the average duration of videos are very short, e.g., 15 seconds for MSRVTT-QA (Xu et al., 2017), 10 seconds for MSVD-QA (Xu et al., 2017), and 3 seconds for TGIF-QA (Jang et al., 2019). The collection and annotation of long video dataset is highly desired. We hope the research and industry committee work together to contribute large datasets of long videos to inspire more investigation in future.
4.3 Evaluation Metrics
In this paper, we follow the metrics of accuracy and average score as proposed by Maaz et al. (2023) for performance evaluation, where we use ChatGPT (chatgpt35-turbo) to assist in judging the correctness of model predictions. ChatGPT accepts questions, groundtruth answers, and the model predictions as input. For each question answer pair, ChatGPT gives a binary judgment of “yes” or “no” to identify whether the predicted answer is correct or not, for accuracy evaluation. Moreover, a score of 0-5 is also given by ChatGPT to indicate how similar the prediction is to the answer. 0 represents the lowest score and 5 represents the highest score.
4.4 Comparison with Other Video Language Models
@@
Table 2: Comparison with the other video-language models, including Video-LLaMA and the baseline method Video-ChatGPT. We report the accuracy (%)/ average score.
Model WildQA QaEgo4D lifeQA Social-IQ 2.0
Video-LLaMA (Zhang et al., 2023a) 63.19/3.18 35.35/1.94 35.75/2.32 55.78/2.90
Video-ChatGPT (Maaz et al., 2023) 58.00/3.30 29.74/2.43 33.87/2.55 57.73/3.26
R-VLM (Ours) 64.82/3.39 32.51/2.45 38.71/2.61 63.65/3.40
We compare our final scheme R-VLM (Retrieval based video language model) with the baseline method Video-ChatGPT (Maaz et al., 2023) on four unseen datasets. We use the same training dataset as Video-ChatGPT. Both Video-ChatGPT and our R-VLM  are built based on LLaVA, with similar model size. Table 2 shows the results. Our R-VLM  outperforms Video-ChatGPT significantly by 6.8%, 2.8%, 4.8%, 6.0% in accuracy on the WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0 dataset, respectively. Ours consistently achieves higher average score. Note that the number of vision tokens of our R-VLM  is comparable to that of Video-ChatGPT (i.e., 340 vs. 356), making a fair comparison. This demonstrates the effectiveness of our retrieval-based design. Our chunk-wise retrieval design facilitates the exploration of the most informative vision tokens and preservation of necessary vision details for QA.
We also evaluate the performance of other video language model, e.g., Video-LLaMA (Zhang et al., 2023a), by testing their model directly. As shown in Table 2, our R-VLM  achieves the best scores on all the four datasets. We found Video-LLaMA is prone to give detailed description of the entire video, whereas the answer is usually question-irrelevant. Fig. 2(a) shows some typical examples on the QaEgo4D dataset, where the answer includes much information while the desired answer is submerge in the overall answer. The accuracy metric is not perfect and trades such answer as correct. In contrast, our model provides more concise and accurate answers.
4.5 Ablation Studies
@@
Table 3: Comparison of different methods and chunk selection strategies. All these models are trained using the same video instruction data. R-VLM denotes our final scheme with learnable retrieval. R-VLM w/ Uni. denotes uniform sampling of $K$ chunks in our framework instead of retrieval-based sampling. R-VLM w/ CLIP M. denotes that we use the final CLIP class token feature of vision and text for matching in our framework, without learnable parameters for retrieval. We report the accuracy (%) / average score.
Dataset Video-ChatGPT R-VLM w/ Uni. R-VLM w/ CLIP M. R-VLM
WildQA 58.00/3.30 61.23/3.36 60.31/3.27 64.82/3.39
QaEgo4D 29.74/2.43 31.57/2.44 31.52/2.43 32.51/2.45
lifeQA 33.87/2.55 36.56/2.56 31.45/2.42 38.71/2.61
Social-IQ 2.0 57.73/3.26 57.96/3.24 61.17/3.28 63.65/3.40
@@
In this section, we study and validate the effectiveness of our retrieval designs, chunk-wise design and the soft matching loss, respectively.
@@
Effectiveness of Retrieval for Chunk Selection. Under our framework, we compare our retrieval mechanism and the uniform sampling strategy for the selection of $K$=5 chunks as context input to the LLM. For the uniform sampling setting which we name as R-VLM w/ Uni., the model selects $K$ from $N$ video chunks uniformly instead of based on question guided retrieval. Table  3 shows that our final scheme R-VLM  with learnable retrieval-based strategy outperforms R-VLM w/ Uni. by 3.6%, 0.9%, 2.2%, and 5.7% on WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0, respectively. For a long video, the video chunks relevant to the question usually account for a small portion of the entire video. It is difficult to hit these question-related chunks with uniform sampling. A large language model may can not correctly answer a question when it accepts video chunks not correlated with the question. In contrast, our learnable retrieval learns to select the chunks most relevant to the question, providing more reliable and less redundancy information to the LLM for effective inference. Fig. 2 which visualizes the selected chunks also validate this. Our retrieval based model is interpretable, providing where the model is based on to get the answer.
@@
Learnable Retrieval vs. Off-the-shelf CLIP based Retrieval. The CLIP image encoder and text encoder are pre-trained to achieve vision and text caption alignment through contrastive learning over the last layer class token features (Radford et al., 2021). One may wonder how about the performance when we use the CLIP class token features for question and chunk feature matching under our framework. To answer this question, we design a scheme R-VLM w/ CLIP M. for comparison. For a chunk, we use the averaged class token feature of the last layer of the CLIP image encoder (CLIP ViT-L/14) as the vision chunk feature, and the class token feature of the last layer of the CLIP text encoder as the question feature for matching. There is no learnable parameters for the retrieval. From Table 3, we can see that our R-VLM  with learnable retrieval consistently outperforms R-VLM w/ CLIP M., upto 7.3% in accuracy on the lifeQA dataset. That may because the CLIP matching is originally designed for image and caption matching, which is not robust to the image and question matching. Thanks to the adaptation of the text encoder through a learnable MLP block, our question guided retrieval can better identify the relevant video chunks.
Effectiveness of Chunk-wise Design. Video-ChatGPT performs video level spatial temporal pooling to obtain 356 vision tokens. Such global pooling would result in loss of details, especially when the question related video segments take a small portion of the entire video. In our design, we perform chunk level spatial temporal pooling to preserve more information about each chunk. Table 3 shows that when we uniform sample the chunks to have 340 vision tokens, R-VLM w/Uni obviously outperforms Video-ChatGPT, demonstrating the effectiveness of our chunk-wise design.
Table 4: Influence of soft matching (SM) loss, evaluated by the accuracy (%) / average score.
Model WildQA QaEgo4D lifeQA Social-IQ 2.0
R-VLM w/o SM 59.94/3.28 31.12/2.36 36.29/2.47 57.22/3.17
R-VLM 64.82/3.39 32.51/2.43 38.71/2.61 63.65/3.40
@@
Influence of Soft Matching (SM) Loss. In order to regularize the learning of the MLP block for better retrieval, we introduce SM loss. Table 4 shows the comparison of our framework without the SM loss (R-VLM w/o SM) and that with SM loss (R-VLM). We can see that incorporating the SM loss significantly improves the performance. Our final scheme with the SM loss i.e., R-VLM  outperforms that without SM Loss. SM Loss uses the cosine similarity between text embedding and video tokens as weight to re-weight the video tokens to obtain new video tokens. Then maximize the similarity between the new video tokens and the text embedding. This facilitates our learnable retrieval layer to find the video clips most similar to the questions.
Influence of the Hyperparameter $K$. In general, when $K$ is too small, it may lead to a loss of information necessary to answer the question. When $K$ is too large, interference may be introduced, confusing the LLM. We found $K$=5 presents a good trade-off on most datasets and set $K$ as 5. More details can be found in our Appendix. We leave the adaptive design of $K$ as the future work.
4.6 Visualization Analysis
We visualize two examples from the QAEgo4D and WildQA datasets in Fig. 2 , with the following information. 1) The first row shows the video chunk samples by uniformly selecting 5 video chunks. 2) The second row shows the retrieved 5 chunks (ordered by time order) in our R-VLM. We mark the groudtruth chunks by red box. 3) We show the learned similarity score (see Eq. (2)) curve based on which the top $K$ chunks are selected. The horizontal axis represents the identity of chunk and the vertical axis denotes the similarity score of that chunk. The groundtruth chunks and our retrieved chunks are also marked. 4) The question and answers from different models: R-VLM, R-VLM w/Uni., Video-ChatGPT, and Video-LLaMA, respectively.
For the QAEgo4D dataset, the average duration is about 8 minutes (120 video chunks), whereas the duration of the groundtruth segments is located 2% of the total video duration in average. It is difficult to hit groundtruth video segments by uniformly sampling $K(=5)$ chunks among 120 video chunks. In Fig. 1(a), we can see that the fragments selected by uniform sampling are in general irrelevant to the problem, which is not conducive to the LLM reasoning. Our learnable retrieval can accurately find the segments where the answer lies in. Feeding the correct chunks to the LLM makes it possible to obtain the correct answer to the question.
For the WildQA dataset, groundtruth segments usually locate at different locations in the video and last for a period of time. Although uniform sampling sometimes hits a certain groundtruth segments, there is no guarantee. In contrast, our learned retrieval can correctly hit the segments where the groundtruth segments are located on. In this way, the LLM can better understand the video and answer some detailed questions (refer to “vegetation types” in Fig. 1(b)).
The predicted answers of our R-VLM  is more accurate than Video-ChatGPT, and Video-LLaMA. More visualization can be found in our Supplementary.
Besides the successful cases, we present some failed cases in our Appendix (see Fig.5). There are two main cases of failure. One is that the retrieval did not select the correct video chunks. The other is that the retrieval correctly identified the correct video chunks, but the answer was wrong. For the later cases, we think more powerful vision feature extractor and LLMs would alleviate the problem.
Figure 2: Visualization of video QA examples from (a) QAEgo4D and (b) WildQA. (a)The kitchen towel in question does not appear in the uniformly sampled video chunks. The second chunks selected by our model contain kitchen towel. Our answer states that the tower is hunging on a hook on the wall. Video-LLaMA answers incorrectly, where the towel does not appear in the first frame of the video, and it is not be placed on the countertop in front of a cutting board. (b)In this video, two clips show vegetation and the remaining clips show mountains, rivers, etc. Uniform sampling mainly obtains segments such as mountains and rivers rather than segments with vegetation. Therefore, only the terrain was answered, without giving vegetation types. In contrast, our retrieved chunks contain video clips of vegetation. Thus the types of vegetation are predicted correctly: trees, bushes, forest. Video-ChatGPT gives a global description and does not answer specific vegetation types.
5 Conclusion
The comprehension of long videos using LLMs remains an under-explored area. There are two main challenges associated with comprehending long videos. 1) Long videos generally lead to abundant vision tokens, which increase computational cost for LLM inference. 2) Global aggregation of vision tokens inevitably results in the loss of vision details especially when the question relevant video chunks take only a small portion of the entire video. Moreover, question irrelevant chunks introduce interference. In this work, we address these issues by introducing a simple yet effective retrieval-based video language model (R-VLM) for long-video QA. Specifically, given a question (query) and a long video, our model identifies and selects the most question-relevant $K$ video chunks and uses their associated visual tokens to serve as context for the LLM inference. This effectively reduces the number of video tokens, preserves the most informative information, eliminates noise interference, and thus enhances system performance. Our experimental results demonstrate the effectiveness of our designs for comprehending long videos.
Acknowledgements
We thank Zhizheng Zhang, Xiaoyi Zhang, Zhipeng Huang, and Tao Yang for their valuable discussion.
Appendix
Appendix A Ablation study on Influence of the Hyperparameter $K$
@@
We study the influence of $K$. In general, when $K$ is too small, it may lead to a loss of information necessary to answer the question. When $K$ is too large, interference may be introduced, confusing the LLM. Table 5 shows the results of using different $K$. We found that as $K$ gradually increases from 1 to 5, the average performance increases. When $K$ increases from 5 to 7, the performance decreases. We found $K$=5 presents a good trade-off on most datasets, even though there are slight differences on different datasets. We leave the adaptive design of $K$ as the future work.
@@
Table 5: Ablation study on the influence of $K$, evaluated in terms of accuracy (%)/score. We use bold to mark the best performance and underline to mark the second-best performance.
@@
Appendix B More visualization results

### v2

@@ III-C End-to-End Optimization
We train the network using an end-to-end optimization approach with the video instruction data. The image encoder, text encoder, and the LLM are frozen during the training. Only the parameters of the MLP block and projector are optimized. For a video-text pair, we introduce soft matching (SM) loss $\mathcal{L}_{i}^{\text{SM}}$ to regularize the similarity learning as
@@
$\mathcal{L}_{i}^{\text{SM}}=-\frac{\mathbf{q}_{i}\cdot\bar{\mathbf{v}}_{i}}{\lVert\mathbf{q}_{i}\rVert\lVert\bar{\mathbf{v}}_{i}\rVert},\quad\text{where}\>\bar{\mathbf{v}}_{i}=\frac{\sum_{j=1}^{L_{i}}e^{s_{i}^{j}}\mathbf{v}_{i}^{j}}{\sum_{j=1}^{L_{i}}e^{s_{i}^{j}}},$ (3)
where $\bar{\mathbf{v}}_{i}$ is a weighted combination of the $\mathcal{L}_{i}$ chunk features. For a weight $e^{s_{i}^{j}}$, $e$ denotes the base of the natural logarithm, $s_{i}^{j}$ denotes the similarity score between the query and the $j^{th}$ chunk features (see Eq. (2)) of the video $V_{i}$. $\mathbf{v}_{i}^{j}$ denotes the averaged token feature of the $j^{th}$ chunk, i.e., $\mathbf{v}_{i}^{j}=\frac{1}{N+M}\sum_{n=1}^{N+M}\mathbf{u}_{i,n}^{j}$. The soft matching loss aims to align the aggregated visual feature with the question feature by adjusting weights. Through the optimization of the MLP block, high weights are assigned to video chunks closely related to the question and low weights to less relevant chunks, thus improving the selection of relevant chunks for accurate question answering.
The overall loss is
@@
where $\mathcal{L}^{pred}_{i}$ denotes the cross-entropy loss between the generated prediction and groundtruth answer (a.k.a. LLM loss). $\lambda$ is a hyper-parameter (we set to 10) that balances the contribution of the regularization term (see Table X for the ablation).
IV Experiments
IV-A Implementation Details
@@
Following Video-ChatGPT [1], we use LLaVA [17] as our base model. We utilize the pre-trained CLIP ViT-L/14 [20] as our image encoder and extract the feature from the second-to-last layer as the $h\times w$ visual tokens of a frame. We use the fine-tuned Vicuna (7B) from LLaVA as our LLM. For the text encoder, we use the pre-trained CLIP ViT-L/14 text encoder and extract the class token feature of the penultimate layer. Each of the two fully connected layers within the MLP block is equipped with 1024 neurons. The projector has 4096 neurons.
@@
We only fine-tune the MLP block and the projector while keeping the image encoder, the text encoder and the LLM frozen. We fine-tune the model for 3 epochs using video instruction data, with a learning rate of 2e-5 and a batch size of 40. Training our model takes about 24h on an A100 80GB GPU. We set $K$ to 5, resulting in $5\times(64+4)=340$ visual tokens as the input to the LLM. This is comparable to the number of visual tokens used in Video-ChatGPT (356 visual tokens).
For VideoChat-GPT, only the projector of 4.2M parameters is trained. In our framework, besides the projector, we added a small trainable MLP block of 1.8M parameters. Our total number of trainable parameters is 6.0M, which is very small considering the total model size is 6.7B.
IV-B Datasets and Evaluation
@@
TABLE I: Information of the evaluation datasets, including the number of videos, the number of QA pairs, and the average video duration.
Dataset #Video #QA Duration (sec.)
Activitynet-QA 800 8000 114
EgoSchema 5031 5031 180
Next-QA 570 4996 44
IntentQA 567 2134 47
WildQA 261 652 71
QaEgo4D 166 1854 495
lifeQA 49 372 74
Social-IQ 2.0 144 876 60
@@
We use Video Instruction Data collected by Maaz et al.[1] for video instruction tuning. This dataset contains about 100k question and answer pairs based on the Activitynet-QA dataset (average duration 180 seconds), including various types of questions.
We evaluate the generalization (zero shot) performance of our framework on long video QA datasets Activitynet-QA [31], EgoSchema [32], NExt-QA[33], and IntentQA [34]. To widely evaluate the effectiveness of our model, besides the four popular datasets, we also tested on other long video QA datasets WildQA [35], QaEgo4D [36], lifeQA [37], and Social-IQ 2.0 [38]. Table I shows specific information about each dataset. We perform ablation studies on Activitynet-QA, WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0. Note that the other widely used datasets MSVD-QA [39], MSRVTT-QA [40] for video QA are not suitable, which have average duration of only 10 seconds and 15 seconds, being too short for evaluation.
We assess model performance using accuracy and a 0-5 similarity score, respectively, as per Video-ChatGPT [1], with ChatGPT (gpt-3.5-turbo) evaluating the correctness of predictions through binary and numerical judgments. See Supplementary Section B-A for more information.
TABLE II: Performance comparison with the state-of-the-art methods for video QA on Activitynet-QA.
Model Acc./Score
Video-LLaMA [4] 12.4/1.1
FrozenBiLM [41] 24.7/-
VideoChat [3] 26.5/2.2
LLaMA-Adapter [42] 34.2/2.7
Elysium [43] 43.4/2.9
Video-LLaVA [44] 45.3/3.3
MovieChat [25] 45.7/3.1
BT-Adapter [45] 46.1/3.2
Chat-UniVi [46] 46.1/3.3
MiniGPT4-video(7B) [47] 46.3/-
LLaMA-VID(7B) [26] 47.4/3.3
Video-ChatGPT [1] 47.8/3.3
VideoChat2 [48] 49.1/3.3
R-VLM (Ours) 49.7/3.3
TABLE III: Performance comparison with the state-of-the-art methods for video QA on EgoSchema fullset.
Model Acc.
SeViLA [24] 22.7
FrozenBiLM [41] 26.9
mPLUG-Owl [22] 31.1
Intervideo [49] 32.1
TimeChat [50] 33.0
LLoVi [51] 33.5
Vamos(13B) [52] 36.7
R-VLM (Ours) 40.2
TABLE IV: Performance comparison with the state-of-the-art methods for video QA on NExt-QA.footnotemark:
Model Des. Tem. Cau. Acc.
Just-Ask [53] 36.0 31.8 30.4 38.4
CLIP [20] 57.0 38.1 43.6 43.9
InternVideo [49] 65.1 43.4 48.0 49.1
Mistral(7B) [54] - - - 51.1
VFC [55] 64.1 45.4 51.6 51.5
LLoVi(7B) [51] 63.2 55.6 47.9 54.3
MVU(13B) [56] 64.1 55.4 48.1 55.2
Video-ChatGPT [1] 62.7 48.8 58.8 56.2
R-VLM (Ours) 63.1 52.0 59.9 58.7
TABLE V: Performance comparison with the state-of-the-art methods for video QA on IntentQA dataset.
Model Acc.
Random 20.0
HGQA [57] 47.7
VGT [58] 51.3
BlindGPT [59] 51.6
Mistral(7B) [54] 50.4
LLoVi(7B) [51] 53.6
CaVIR [34] 57.6
R-VLM (Ours) 57.7
@@
TABLE VI: Comparison with the other video-language models, including Video-LLaMA [23] and the baseline method Video-ChatGPT [1]. We report the accuracy (%)/ average score.
Dataset Video-LLaMA Video-ChatGPT R-VLM (Ours)
WildQA 63.19/3.18 58.00/3.30 64.82/3.39
QaEgo4D 35.35/1.94 29.74/2.43 32.51/2.45
lifeQA 35.75/2.32 33.87/2.55 38.71/2.61
Social-IQ 2.0 55.78/2.90 57.73/3.26 63.65/3.40
IV-C Comparison with Other Models
We compare our final scheme R-VLM (Retrieval based video-language model) with the baseline method Video-ChatGPT [1] and the other state-of-the-art methods. We treat Video-ChatGPT as our baseline method, where we use the same training dataset, LLM of LLaMA-7B, and visual encoder CLIP-L as Video-ChatGPT.
Table II, III, IV, and Table VI show the performance comparison on Activitynet-QA, EgoSchema, NExt-QA, and IntentQA, respectively. We can see that our model achieves the state-of-the-art performance. On Activitynet-QA, our R-VLM outperforms LLaMA-VID-7B by 2.3% in accuracy. Our R-VLM achieves competitive performance with VideoChat2, even though VideoChat2 uses much more training data than ours (1.1 million vs. our 100K video-text pairs).
On the WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0 datasets, Table VI shows the comparison with the baseline method Video-ChatGPT, and the method of Video-LLaMA [23]. R-VLM outperforms Video-ChatGPT significantly by 6.8%, 2.8%, 4.8%, 6.0% in accuracy, respectively. Ours consistently achieves the best average score. Note that the number of visual tokens of our R-VLM is comparable to that of Video-ChatGPT (i.e., 340 vs. 356), making a fair comparison. This demonstrates the effectiveness of our learnable retrieval-based design, which facilitates the exploration of the most informative visual tokens and preservation of necessary vision details for QA. We found that when evaluating Video-LLaMA, the metric of accuracy is not as reliable as the score. Video-LLaMA is prone to give detailed description of the entire video, whereas the answer is usually not question-specific. The accuracy metric treats such answer as correct, while the score rating is relatively reasonable. Fig. 5 in Supplementary Material B-E shows a typical example on QaEgo4D, where the answer includes much irrelevant (or incorrect) information while the answer is submerged in the overall answer. In contrast, our model provides more concise and accurate answers. This explains why our accuracy is lower than Video-LLaMA on QaEgo4D.
footnotetext: Sevilla [24] achieves an average accuracy of 73.8%, where the model is trained on NExT-QA. In contrast, our model has not seen Next-QA during training. VideoChat2 [48] achieves an average accuracy of 68.6%. It uses much more data for training (1.9M samples vs. ours using 100K).
@@ IV-C Comparison with Other Models
TABLE VII: Comparison of different methods and ablation studies. All these models are trained using the same video instruction data. R-VLM denotes our final scheme with learnable retrieval. Video-ChatGPT is our baseline method. R-VLM w/ Uni. denotes uniform sampling of $K$ chunks in our framework instead of retrieval-based sampling. R-VLM w/ CLIP M. denotes that we use the final CLIP class token features of vision and text for matching in our framework, without incoperating learnable parameters for retrieval. R-VLM w/o SM denotes that we do not use the proposed soft matching (SM) loss. R-VLM w/o G. denotes that the gradient back-propagation from the LLM loss $\mathcal{L}^{pred}_{i}$ to the learnable MLP block $\psi$ is disabled. We report the accuracy (%) / average score.
Dataset Video-ChatGPT R-VLM w/ Uni. R-VLM w/ CLIP M. R-VLM w/o SM R-VLM w/o G. R-VLM
ActivityNet-QA 47.81/3.27 49.34/3.32 48.86/3.30 48.39/3.32 48.30/3.31 49.65/3.34
WildQA 58.00/3.30 61.23/3.36 60.31/3.27 59.94/3.28 62.27/3.32 64.82/3.39
QaEgo4D 29.74/2.43 31.57/2.44 31.52/2.43 31.12/2.36 31.66/2.46 32.51/2.45
lifeQA 33.87/2.55 36.56/2.56 31.45/2.42 36.29/2.47 37.09/2.60 38.71/2.61
Social-IQ 2.0 57.73/3.26 57.96/3.24 61.17/3.28 57.22/3.17 62.43/3.36 63.65/3.40
IV-D Ablation Studies
@@
In this section, we study and validate the effectiveness of our retrieval designs, the soft matching loss, chunk-wise design, the influence of hyperparameter $K$, and the influence of the time duration of a chunk, respectively.
@@
Effectiveness of Retrieval for Chunk Selection Within our framework, we compare our retrieval mechanism and the uniform sampling strategy for the selection of $K=5$ chunks as context input to the LLM. For the uniform sampling setting which we name as R-VLM w/ Uni., the model selects $K$ from $N$ video chunks uniformly instead of based on question guided retrieval. Table  VII shows that our final scheme R-VLM with learnable retrieval-based strategy outperforms R-VLM w/ Uni. by 3.6%, 0.9%, 2.2%, and 5.7% on WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0, respectively. For a long video, the video chunks relevant to the question usually account for a small portion of the entire video. It is difficult to hit these question-related chunks with uniform sampling. A large language model may not correctly answer a question when it accepts video chunks not correlated with the question. In contrast, our learnable retrieval learns to select the chunks most relevant to the question, providing more reliable and less redundant information to the LLM for effective inference. Visualizations of the selected chunks in Fig. 2 also validate this and show that our retrieval-based model is interpretable, indicating from where and why the model produces the answer.
@@
Effectiveness of Learnable Retrieval vs. Off-the-shelf CLIP based Retrieval We propose the learnable retrieval by introducing the MLP block. One may wonder the performance of using off-the-shelf CLIP class token features to match question and chunk features inside our framework. To answer this, we design a scheme R-VLM w/ CLIP M. for comparison. For a chunk, we use the averaged class token feature of the last layer of the CLIP image encoder as the vision chunk feature, and the class token feature of the last layer of the CLIP text encoder as the question feature for matching. There is no learnable parameters for the retrieval. Table VII shows that our R-VLM with learnable retrieval consistently outperforms R-VLM w/ CLIP M., upto 7.3% in accuracy on the lifeQA dataset. That may be because the CLIP matching is originally designed for image and caption matching, which is not robust to the image and question matching. Thanks to the adaptation of the text encoder through a learnable MLP block, our question guided retrieval can better identify the relevant video chunks.
@@
Influence of Soft Matching (SM) Loss In order to regularize the learning of the MLP block for better retrieval, we introduce SM loss. Table VII shows the comparison of our framework without the SM loss (R-VLM w/o SM) and that with SM loss (R-VLM). We can see that incorporating the SM loss significantly improves the performance. Our final scheme with the SM loss i.e., R-VLM outperforms that without SM Loss. SM Loss uses the cosine similarity between text embedding and video tokens as weight to re-weight the video tokens to obtain new video tokens. Then maximize the similarity between the new video tokens and the text embedding. This facilitates our learnable retrieval layer to find the video clips most similar to the questions.
Influence of LLM Loss to Retrieval The instruction tuning with the LLM loss ($\mathcal{L}^{pred}_{i}$) back-propagates the gradient also to the learnable MLP block $\psi$, which facilitates the selection of right chunks for retrieval. When we disable the back-propagation to $\psi$, we have scheme R-VLM w/o G.. As shown in Table VII, our final scheme R-VLM that allows the back-propagation obviously outperforms R-VLM w/o G. by 2.6%, 0.9%, 1.6%, and 1.2% on WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0.
Effectiveness of Chunk-wise Design Video-ChatGPT performs video level spatial temporal pooling to obtain 356 visual tokens. Such global pooling would result in loss of details, especially when the question-related video segments take a small portion of the entire video. In our design, we perform chunk level spatial temporal pooling to preserve more information about each chunk. Table VII shows that when we uniformly sample the chunks to have 340 visual tokens, R-VLM w/Uni obviously outperforms Video-ChatGPT, demonstrating the effectiveness of our chunk-wise design.
@@
Influence of Hyperparameter $K$ In general, when $K$ is too small, it may lead to a loss of information necessary to answer the question. When $K$ is too large, interference may be introduced, confusing the LLM. Table VIII shows the results of using different $K$. We found that as $K$ gradually increases from 1 to 5, the average performance increases. When $K$ increases from 5 to 7, the performance decreases. We found $K=5$ presents a good trade-off on most datasets, even though there are slight differences on different datasets. We leave the adaptive design of $K$ as future work.
@@
TABLE VIII: Ablation study on the influence of $K$, evaluated in terms of accuracy (%)/score. We use bold to mark the best performance and underline to mark the second-best performance.
@@
Influence of the Time Duration of a Chunk In considering the memory limitation and video redundancy, we sample at a low frame rate (1 fps). To allow a chunk to contain dynamic temporal information, we use 4 seconds that are expected to contain temporal dynamics as the chunk unit. We study the influence of the time duration of a chunk by comparing the use of 4 frames (4 seconds) and 8 frames (8 seconds) as a chunk at the same sample rate of 1 fps. Table IX shows the performance comparison. We can see that the performance of using 4 frames/chunk is better than that using 8 frames/chunk. Firstly, 4 frames/chunk allows identification of shorter temporal segments. Secondly, it preserves more spatial details than 8 seconds per chunk since an 8-second chunk pools features over a longer period. Both the two settings outperform the baseline method Video-ChatGPT.
TABLE IX: Influence of the time duration of a chunk, evaluated in terms of accuracy (%)/score. 8f/c denotes 8frames/chunk.
Video-ChatGPT R-VLM(8f/c) R-VLM(4f/c)
ActivityNet-QA 47.81/3.27 49.24/3.31 49.65/3.34
WildQA 58.00/3.30 62.58/3.37 64.82/3.39
QaEgo4D 29.74/2.43 32.36/2.43 32.51/2.45
lifeQA 33.87/2.55 36.83/2.63 38.71/2.61
Social-IQ 2.0 57.73/3.26 62.20/3.33 63.65/3.40
Influence of the Hyperparameter $\lambda$ We determine $\lambda$ based on our previous experiences, where we usually can obtain satisfactory results by setting different losses at the same magnitude. This is a simple way to ensure that different losses are considered in optimization. We found when we set $\lambda$ as 10, the two loss terms are at the similar magnitude. We also conducted ablation study on the choose of $\lambda$ by setting it to 1, 10, and 100. Table X shows the comparison. We can see that $\lambda=10$ provides he best performance on average, slightly better than that of $\lambda=1$.
TABLE X: Ablation study on $\lambda$.
$\lambda$=1 $\lambda$=10 $\lambda$=100
ActivityNet-QA 49.70/3.32 49.65/3.34 49.06/3.33
WildQA 64.42/3.41 64.82/3.39 63.34/3.35
QaEgo4D 34.57/2.50 32.51/2.45 32.85/2.48
lifeQA 36.49/2.56 38.71/2.61 36.93/2.67
Social_IQ 2.0 62.77/3.38 63.65/3.40 60.71/3.32
Average 49.59/3.03 49.87/3.04 48.58/3.03
Figure 2: Visualization of video QA examples from QAEgo4D. The kitchen towel related to the question does not appear in the uniformly sampled video chunks. The second chunk selected by our model contains kitchen towel. Our answer states that the towel is hunging on a hook on the wall. Video-LLaMA answers incorrectly, where the towel does not appear in the first frame of the video, and it is not be placed on the countertop in front of a cutting board. Note that we use a sampled frame in a chunk to illustrate the chunk in the first two rows.
IV-E Visualization Analysis
We visualize an example from QAEgo4D in Fig. 2. More examples from QAEgo4D and WildQA are in Fig. 5, Fig. 5 in Supplementary Material B-E, with the following information.